### v2

@@ Abstract
As the economic and environmental costs of training and deploying large vision or language models increase dramatically, analog in-memory computing (AIMC) emerges as a promising energy-efficient solution. However, the training perspective, especially its training dynamic, is underexplored. In AIMC hardware, the trainable weights are represented by the conductance of resistive elements and updated using consecutive electrical pulses. Among all the physical properties of resistive elements, the response to the pulses directly affects the training dynamics. This paper first provides a theoretical foundation for gradient-based training on AIMC hardware and studies the impact of response functions. We demonstrate that noisy update and asymmetric response functions negatively impact Analog SGD by imposing an implicit penalty term on the objective. To overcome the issue, Tiki-Taka, a residual learning algorithm, converges exactly to a critical point by optimizing a main array and a residual array bilevelly. The conclusion is supported by simulations validating our theoretical insights.
@@ 1 Introduction
The remarkable success of large vision and language models is underpinned by advances in modern hardware accelerators, such as GPU, TPU (Jouppi et al., 2023), NPU (Esmaeilzadeh et al., 2012), and NorthPole chip (Modha et al., 2023). However, the computational demands of training these models are staggering. For instance, training LLaMA (Touvron et al., 2023) cost \$2.4 million, while training GPT-3 (Brown et al., 2020) required \$4.6 million, highlighting the urgent need for more efficient computing hardware. Current mainstream hardware relies on the Von Neumann architecture, where the physical separation of memory and processing units creates a bottleneck due to frequent and costly data movement between them.
In this context, the industry has turned its attention to analog in-memory computing (AIMC) accelerators based on resistive crossbar arrays (Chen, 2013; Sebastian et al., 2020; Haensch et al., 2019; Sze et al., 2017), which excel at accelerating the ubiquitous and computationally intensive MVM operations. In AIMC hardware, the weights (matrices) are represented by the conductance states of the resistive elements in analog crossbar arrays (Burr et al., 2017; Yang et al., 2013), while the input and output of MVM are analog signals like voltage and current. Leveraging Kirchhoff’s and Ohm’s laws, AIMC hardware achieves 10$\times$-10,000$\times$ energy efficiency than GPU (Jain et al., 2019; Cosemans et al., 2019; Papistas et al., 2021) in the model inference.
Despite its high efficiency, analog training is considerably more challenging than inference since it involves frequent weight updates. Unlike digital hardware, where the weight increment can be applied to the original weight in the memory cell, the weights in AIMC hardware are changed by the so-called pulse update. When receiving electrical pulses from its peripheral circuits, the resistive elements change their conductance as a response according to the pulse polarity (Gokmen & Vlasov, 2016). However, the conductance of resistive elements increases and decreases following different response curves, leading to asymmetric update.
@@
Figure 1: The weight’s response curve. Positive and negative pulses are sent continuously on the left and right half, respectively. One pulse is sent to the resistive element per cycle. Given $w$, the weight becomes $w^{+}$ or $w^{-}$ after one positive and negative pulse, respectively. The response factors $q_{+}(w)$ and $q_{-}(w)$ are approximately the slope of the curve at $w$, and $\Delta w_{\min}$ is the response granularity. (Left) Ideal response functions $q_{+}(w)=q_{-}(w)$. Every point is symmetric points. (Right) Asymmetric response functions $q_{+}(w)\neq q_{-}(w)$ almost everywhere expect for the symmetric point $w^{\diamond}$.
Ideally, the response to each pulse should remain unchanged regardless of time or conductance state variation, which enables concise control of the weight update. A series of works seek various resistive elements that have near-constant or at least symmetric responses. The leading candidates currently include PCM (Burr et al., 2016; Le Gallo & Sebastian, 2020), ReRAM (Jang et al., 2014; 2015; Stecconi et al., 2024), ECRAM (Tang et al., 2018; Onen et al., 2022), to name a few. Currently, there is still debate regarding the most suitable resistive element for AIMC hardware. Anticipating to cover the wide ranges of resistive element candidates with various response functions, this paper studies generic response under a set of mild assumptions. The key concept is the response function of resistive elements. We are interested in how the resistive elements determine the response functions and how the response functions affect the convergence properties of analog training algorithms.
1.1 Analog training and its challenges
Pulse update and asymmetric issue. In AIMC hardware, the weights of a model are stored in crossbar arrays by the conductance states of resistive elements. To modify the conductance (namely the model weights), a series of electrical pulses need to be sent to resistive elements in consecutive pulse cycles; see Figure 1. Receiving a pulse at each pulse cycle, the conductance is updated by a small amount around $\Delta w_{\min}$, which is called response granularity. However, resistive elements’ response varies from the positive and negative directions, reflected by different slopes in their response curves. The slopes, $q_{+}(w)$ and $q_{-}(w)$, is called response functions in this paper. All $\Delta w_{\min}$, $q_{+}(w)$, and $q_{-}(w)$ are element-dependence parameters or functions, which are set before the training process and hence kept fixed during the training. The asymmetric update leads to an asymmetric response curve and an imperfect training dynamic.
@@ 1.1 Analog training and its challenges
Supported by pulse update, the gradient-based training algorithms are used to optimize the weights. Consider a standard training problem with a model parameterized by $W\in{\mathbb{R}}^{D}$
@@
where $f(\cdot):{\mathbb{R}}^{D}\to{\mathbb{R}}$ is the objective function and $\xi$ is a random variable sampled from an unknown distribution.
Gradient-based training implemented by analog update. Similar to stochastic gradient descent (SGD) in digital training (Digital SGD), the gradient-based training algorithm on AIMC hardware, Analog SGD, updates the weights by the gradient of the objective function. Let $\nabla f(W_{k};\xi_{k})$ be the stochastic gradient of $f(W_{k};\xi_{k})$ at $W_{k}$ with the random variable $\xi_{k}$. The SGD updates the weight by $W_{k+1}=W_{k}-\alpha\nabla f(W_{k};\xi_{k})$ with learning rate $\alpha$. Analog SGD implements it by sending a pulse series with length about $|[\Delta W]_{i}|/\Delta w_{\min}$ to the $i$-th element, where $\Delta W=-\alpha\nabla f(W_{k};\xi_{k})$ is the desired update. With each pulse incurring roughly by $\Delta w_{\min}$, $W_{k}$ is updated by $\Delta W$. This process involving multiple pulse cycles to apply desired update $\Delta W$ is called Analog Update.
However, it is observed that Analog SGD suffers from a serious convergence issue due to the asymmetric update. To alleviate this issue, a family of heuristic variants, Tiki-Taka, (Gokmen & Haensch, 2020; Gokmen, 2021; Rasch et al., 2024) is proposed which introduces another crossbar array to accumulate the gradient. Recently, Wu et al. (2024a) provides a theoretical justification for Tiki-Taka and shows that Tiki-Taka outperforms Analog SGD by eliminate the asymptotic error in Analog SGD caused by the asymmetric update. However, their work is limited to a special case of linear response, which are in the form of $q_{+}(w)=1-w/{\tau},q_{-}(w)=1+w/{\tau}$ with hardware-specific parameter $\tau>0$. Given more general $q_{+}(w)$ and $q_{-}(w)$, the convergence of Tiki-Taka does not trivially holds.
Challenges with generic response functions. It is not necessary for Tiki-Taka to outperform Analog SGD on generic resistive elements, even though the response functions are still linear. Consider a more generic linear response setting $q_{+}(w)=(1+c_{\texttt{Lin}})(1-w/{\tau}),q_{-}(w)=(1-c_{\texttt{Lin}})(1+w/{\tau})$ with a parameter $c_{\texttt{Lin}}$, which reduces to the setting in Wu et al. (2024a) when $c_{\texttt{Lin}}=0$. The modification is slight, but it harms the convergence of Tiki-Taka significantly.
Figure 2: Objective comparison of Analog SGD and Tiki-Taka under different parameter $c_{\texttt{Lin}}$. The error plateau in the order $10^{-5}$ comes from the limited response granularity $\Delta w_{\min}=10^{-4}$.
Figure 2 shows the damage from a non-zero $c_{\texttt{Lin}}$ to Tiki-Taka. Consistent with the conclusion in Wu et al. (2024a), Tiki-Taka significantly outperforms Analog SGD when $c_{\texttt{Lin}}=0$. However, when $c_{\texttt{Lin}}$ is perturbed from $0.1$ to $0.3$, Tiki-Taka degrades dramatically and even becomes worse than Analog SGD does, indicating the non-negligible impact of the response functions. Therefore, there is no convergence guarantee for Tiki-Taka on generic responses, even with linear response. It motivates us to study the condition of the response functions for convergence.
@@ 1.2 Main results
Complementing existing empirical studies in analog in-memory computing, this paper aims to build a rigorous foundation of analog training with generic response functions. This paper builds upon the rich study of various resistive elements, including PCM, ReRAM, and so on, where the behavior of the resistive element at each pulse cycle is well studied. However, at the algorithm level, we are more interested in the dynamic of Analog Update, which describes the process of applying a desired update $\Delta W$ to the weight $W$ implemented by sending pulse series with a proper number. Analog Update hides the details on the management of pulse series and enables us to study the convergence behavior easily. It raises a natural question:
Q1) How do the resistive elements determine the dynamic of Analog Update?
This paper finds that the response functions are the key to connecting pulse update and Analog Update. We start by studying the behavior of the resistive element at each pulse cycle. Based on that, we propose an approximated dynamic of Analog Update and analyze the approximation error.
The response functions are the main contributor to the difference between digital and analog training. It motivates us to study a class of benign response functions friendly to the convergence of Tiki-Taka and raises another question:
Q2) What is the impact of generic response functions on analog training?
To answer the question above, we establish the convergence of Analog SGD and Tiki-Taka on generic response functions based on the proposed dynamic.
Our contributions. The contributions of the paper include:
@@ 1. C1)
    Building upon the equation of pulse update, we propose an approximated dynamic of Analog Update which hides non-essential details in the resistive element level while remaining mathematically tractable. Enabled by it, we study the impact of response functions directly without being limited to concrete element candidates.
@@ 2. C2)
    Based on critical physical properties, we focus on a class of benign response functions. Based on that, we show that analog training suffers an implicit penalty. It attracts the weights towards symmetric points and causes inexact convergence of Analog SGD.
@@ 3. C3)
    We demonstrate the method to alleviate the asymmetric update and implicit penalty issues by residual learning, which introduces another sequence, called residual array, with stationary point 00. This method leads to Tiki-Taka heuristically proposed in Gokmen & Haensch (2020). By properly zero-shifting so that the stationary and symmetric points overlap, Tiki-Taka provably converges to a critical point.
2 Dynamic of Non-ideal Analog Update
The primary distinction between digital and analog training is the weight update method. As discussed in Section 1, the weight update in AIMC hardware is implemented by Analog Update, which sends a series of pulses to the resistive elements. This section provides an approximated dynamic of Analog Update which focuses on the essential details in the resistive element level.
Pulse update. Consider the response of one resistive element in one cycle, which involves only one pulse. Given the initial weight $w$, the updated weight increases or decreases by about $\Delta w_{\min}$ depending on the pulse polarity, where $\Delta w_{\min}>0$ is the response granularity determined by elements. The granularity is further scaled by a factor, which varies by the update direction due to the asymmetric update property of resistive elements. The notations $q_{+}(\cdot)$ and $q_{-}(\cdot)$ are used to denote the response functions on positive or negative sides, respectively, to describe the dominating part of the factor. In practice, the analog noise also causes a deviation of the effective factor from the response functions, referred to as cycle variation. It is represented by the magnitude $\sigma_{c}$ times a random variable $\xi_{c}$ with expectation 00 and variance $1$. Taking all of them into account, with $s\in\{+,-\}$ being the update direction, the updated weight after receiving one pulse is ${\tilde{U}}_{q}(w,s)$ where ${\tilde{U}}_{q}(\cdot,\cdot):{\mathbb{R}}\times\{+,-\}\to{\mathbb{R}}$ is the element-dependent update that implements the resistive element, which can be expressed as
${\tilde{U}}_{q}(w,s):=$ $\ w+\Delta w_{\min}\cdot(q_{s}(w)+\sigma_{c}\xi)$ (2)
The typical signal and noise ratio $\sigma_{c}/q_{s}(w)$ is roughly 5%-100% (Gong et al., 2018; Stecconi et al., 2024), varied by the type of resistive elements. Furthermore, the response functions also vary by elements due to the imperfection in fabrication, called element variation (also referred to as device variation in literature Gokmen & Vlasov (2016)).
Equation (2) is a resistive element level equation. Existing work exploring the candidates of resistive elements usually reports the response curves similar to Figure 1, (Gong et al., 2022; Tang et al., 2018; Stecconi et al., 2024). Taking the difference between weights in two consecutive pulse cycles and adopting statistical approaches Gong et al. (2018), all the element-dependent quantities, including $\Delta w_{\min}$, $q_{+}(\cdot)$, $q_{-}(\cdot)$ and $\sigma_{c}$, can be estimated from the response curves of the resistive elements.
Analog update implemented by pulse updates. Even though the update scheme has evolved over the years (Gokmen & Vlasov, 2016; Gokmen et al., 2017), we discuss a simplified version, called Analog Update, to retain the essential properties. To update the weight $w$ by $\Delta w$, a series of pulses are sent, whose bit length (BL) is computed by $\operatorname{BL}:=\left\lceil\frac{|\Delta w|}{\Delta w_{\min}}\right\rceil$. After received $\operatorname{BL}$ pulses, the updated weight $w^{\prime}$ can be expressed as the function composition of (2) by $\operatorname{BL}$ times
Roughly speaking, given an ideal response $q_{+}(w)=q_{-}(w)=1$ and $\sigma_{c}=0$, $\operatorname{BL}$ pulses, with $\Delta w_{\min}$ increment for each individual pulse, incur the weight update $\Delta w$. Since the response granularity $\Delta w_{\min}$ is scaled by the response function $q_{s}(w)$, the expected increment is approximately scaled by $q_{s}(w)$ as well. Accordingly, we propose an approximate dynamic of Analog Update is given by $w^{\prime}\approx U_{q}(w,\Delta w)$, where $U_{q}(w,\Delta w)$ is defined by

### v3

@@ Abstract
As the economic and environmental costs of training and deploying large vision or language models increase dramatically, analog in-memory computing (AIMC) emerges as a promising energy-efficient solution. However, the training perspective, especially its training dynamic, is underexplored. In AIMC hardware, the trainable weights are represented by the conductance of resistive elements and updated using consecutive electrical pulses. While the conductance changes by a constant in response to each pulse, in reality, the change is scaled by asymmetric and non-linear response functions, leading to a non-ideal training dynamic. This paper provides a theoretical foundation for gradient-based training on AIMC hardware with non-ideal response functions. We demonstrate that asymmetric response functions negatively impact Analog SGD by imposing an implicit penalty on the objective. To overcome the issue, we propose residual learning algorithm, which provably converges exactly to a critical point by solving a bilevel optimization problem. We demonstrate that the proposed method can be extended to address other hardware imperfections, such as limited response granularity. As we know, it is the first paper to investigate the impact of a class of generic non-ideal response functions. The conclusion is supported by simulations validating our theoretical insights.
@@ 1 Introduction
The remarkable success of large vision and language models is underpinned by advances in modern hardware accelerators, such as GPU, TPU [1], NPU [2], and NorthPole chip [3]. However, the computational demands of training these models are staggering. For instance, training LLaMA [4] cost \$2.4 million, while training GPT-3 [5] required \$4.6 million, highlighting the urgent need for more efficient computing hardware. Current mainstream hardware relies on the Von Neumann architecture, where the physical separation of memory and processing units creates a bottleneck due to frequent and costly data movement between them.
[…]
@@
Figure 1: The weight’s response curve. Positive and negative pulses are fired continuously on the left and right halves, respectively. One pulse is fired per cycle. Given $w$, the weight becomes $w^{+}$ or $w^{-}$ after one positive and negative pulse, respectively. The response factors $q_{+}(w)$ and $q_{-}(w)$ are approximately the slope of the curve at $w$, and $\Delta w_{\min}$ is the response granularity. (Left) Ideal response functions $q_{+}(w)\equiv q_{-}(w)$. Every point is a symmetric point. (Right) Asymmetric response functions $q_{+}(w)\neq q_{-}(w)$ almost everywhere expect for the symmetric point $w^{\diamond}$.
Despite its high efficiency, analog training is considerably more challenging than inference since it involves frequent weight updates. Unlike digital hardware, where the weight increment can be applied to the original weight in the memory cell, the weights in AIMC hardware are changed by the so-called pulse update. Pulse update. When receiving electrical pulses from its peripheral circuits, the resistive elements change their conductance in response to the pulse polarity [16]. Receiving a pulse at each pulse cycle, the conductance is updated by a small amount around $\Delta w_{\min}$, which is called response granularity. $\Delta w_{\min}$ is scaled by a response functions $q_{+}(w)$ or $q_{-}(w)$ depending on its polarity. Geometrically, $q_{+}(w)$ and $q_{-}(w)$ are the slopes of response curves; see Figure 1. All $\Delta w_{\min}$, $q_{+}(w)$, and $q_{-}(w)$ are element-specific parameters or functions that are set before training and hence remain fixed during training. Typically, $\Delta w_{\min}$ is known while $q_{+}(w)$ and $q_{-}(w)$ are not.
@@
Gradient-based training implemented by analog update. Supported by pulse update, the gradient-based training algorithms are used to optimize the weights. Consider a standard training problem with objective $f(\cdot):{\mathbb{R}}^{D}\to{\mathbb{R}}$ and a model parameterized by $W\in{\mathbb{R}}^{D}$
@@
where $\xi$ is a random data sample. Similar to stochastic gradient descent (SGD) in digital training (Digital SGD), the gradient-based training algorithm on AIMC hardware, Analog SGD, updates the weights by stochastic gradients $\nabla f(W_{k};\xi_{k})$. Digital SGD updates the weight by $W_{k+1}=W_{k}-\alpha\nabla f(W_{k};\xi_{k})$ with learning rate $\alpha$. Given a desired update $\Delta W=-\alpha\nabla f(W_{k};\xi_{k})$, AIMC hardware implements Analog SGD by sending about $|[\Delta W]_{i}|/\Delta w_{\min}$ pulses to the $i$-th element. With each pulse updating $W_{k}]_{i}$ by $\Delta w_{\min}$, $W_{k}]_{i}$ is ultimately updated by about $\Delta W]_{i}$.
Challenges of analog training. Despite its ultra-efficiency, gradient-based training on AIMC hardware is challenging. First, the generic response functions are asymmetric (i.e. $q_{+}(w)\not\equiv q_{-}(w)$), and non-linear [17, 18, 19]. Due to the variation of response functions and conductance states, gradients are scaled by different magnitudes (i.e., $q_{s}([W_{k}]_{i})$ for $s=\pm$) across different coordinates, leading to biased gradients. Furthermore, the response granularity $\Delta w_{\min}$ is a constant. When the gradients or the learning rate decay below $\Delta w_{\min}$, pulse update no longer provides sufficient precision to perform gradient descent [20]. Other imperfections include, but are not limited to, input/output (IO) of MVM operations and analog-digital conversion error [18]. This paper aims to investigate the impact of non-ideal response functions and develop a method to mitigate their negative effects. We also discuss extending the proposed method to deal with other hardware imperfections.
@@ 1.1 Main results
Complementing existing empirical studies in analog in-memory computing, this paper aims to build a rigorous theoretical foundation of analog training. By introducing bias to the gradient, the asymmetric response function plays a central role in differentiating digital and analog training. In contrast, the other non-idealities hinder the training process by causing precision-related issues. Therefore, we approach the problem progressively, beginning with a simplified case that involves only the asymmetric response functions, and extending the proposed methods to more general scenarios.
As a warm-up, building upon the pulse update mechanism, we propose the following discrete-time mathematical model to characterize the trajectory of Analog SGD
$\texttt{Analog~SGD}\qquad W_{k+1}=$ $\ W_{k}-\alpha\nabla f(W_{k};\xi_{k})\odot F(W_{k})-\alpha|\nabla f(W_{k};\xi_{k})|\odot G(W_{k})$ (2)
where $\alpha>0$ is the learning rate and $\xi_{k}$ is the data sample of iteration $k$; $|\cdot|$ and $\odot$ represent the element-wise absolute value and multiplication, respectively; and $F(\cdot)$ and $G(\cdot)$ are hardware-specific matrix which are defined by $q_{+}(\cdot)$ and $q_{-}(\cdot)$. In Section˜2, we will explain the underlying rationale of (2). Compared with the standard Digital SGD, the gradients in (2) are scaled by $F(\cdot)$ and an extra bias term is introduced. Typically, hardware imperfections lead to non-ideal response functions, i.e., $F(\cdot)\not\equiv 1$ and $G(\cdot)\not\equiv 0$. Thus, we ask a natural question that
Q1) What is the impact of non-ideal response functions and how to alleviate it?
Recently, [21] partially answers the question by showing Analog SGD suffers from a convergence issue due to the asymmetric update, and a heuristic algorithm, Tiki-Taka [22, 23, 24], converges exactly by reducing the weight drift. However, their work is limited to a special case of linear response, which are in the form of $q_{+}(w)=1-w/{\tau},q_{-}(w)=1+w/{\tau}$ with hardware-specific parameter $\tau>0$. Given more general $q_{+}(w)$ and $q_{-}(w)$, the convergence of Tiki-Taka does not trivially holds, even though the response functions are still linear.
Figure 2: Comparison of Analog SGD and Tiki-Taka under different parameter $c_{\texttt{Lin}}$. The error plateau in the order $10^{-5}$ comes from the limited response granularity $\Delta w_{\min}=10^{-4}$.
Gap between theory for special linear and generic response. Consider a more generic linear response setting $q_{+}(w)=(1+c_{\texttt{Lin}})(1-w/{\tau}),q_{-}(w)=(1-c_{\texttt{Lin}})(1+w/{\tau})$ with a parameter $c_{\texttt{Lin}}$, which reduces to the setting in [21] when $c_{\texttt{Lin}}=0$. Figure 2 shows the damage from a non-zero $c_{\texttt{Lin}}$ to Tiki-Taka. Consistent with the conclusion in [21], Tiki-Taka significantly outperforms Analog SGD when $c_{\texttt{Lin}}=0$. However, when $c_{\texttt{Lin}}$ is perturbed from $0.1$ to $0.3$, Tiki-Taka degrades dramatically and even becomes worse than Analog SGD does. The modification is slight, but the convergence guarantee in [21] fails, and the convergence of Tiki-Taka is harmed significantly. This counter-example indicates a gap between the theory with special linear response and generic response, and necessitates the study of the analog training with generic response functions and the exploration of exact convergence conditions.
Ignoring other imperfections temporarily, we first analyze the impact of response functions. We show that Analog SGD suffers from asymptotic error due to the mismatch between the algorithmic stationary point and physical symmetric point. Inspired by that, we propose a novel algorithm framework that aligns two points, overcoming the asymmetric issues. Building on that, we endeavor to extend the proposed algorithm to more practical scenarios that involve other imperfections like limited granularity and noisy readings, prompting a second critical question:
Q2) How to extend the framework to address the limited response granularity and noisy IO issues?
To answer this question, we propose two mechanisms to further overcome these two issues.
Our contributions. This paper makes the following contributions:
@@ 1. C1)
    Building upon the equation of pulse update, we propose an approximated discrete-time dynamics of analog update. Enabled by it, we study the impact of response functions directly without being limited to concrete element candidates.
@@ 2. C2)
    Based on that, we show that instead of optimizing $f(\cdot)$, Analog SGD optimizes another penalized objective implicitly. An implicit penalty is introduced by the asymmetric response functions and attracts the weights towards symmetric points. Consequently, Analog SGD can only converge to the optimal point inexactly.
@@ 3. C3)
    We propose a novel Residual Learning algorithm framework to alleviate the asymmetric update and implicit penalty issues. Residual Learning explicitly introduces another residual array, which has a stationary point 0. By properly zero-shifting so that the stationary and symmetric points overlap, Residual Learning provably converges to a critical point.
  4. C4)
    Building on C3), we propose a variant, Residual Learning v2, tailored for more practical training scenarios. We propose introducing a digital buffer to filter out the reading error due to the IO noise. Furthermore, we propose a threshold-based transfer rule to alleviate the instability due to the limited granularity.
1.2 Prior art
AIMC training. Analog training has shown promising early successes with tremendous energy advantage [25, 26]. Among them, on-chip training, which performs forward, backward, and update directly on analog chips [22, 23, 24, 27, 28] is considered to be the most efficient paradigm, but it is more sensitive to hardware imperfections. Sacrificing energy efficiency for robustness, hybrid digital-analog off-chip training is proposed [29, 30, 31, 32], which offloads some computation burden to digital components. This paper focuses on the more challenging on-chip training setting.
Energy-based model and equilibrium propagation. AIMC training leverages back-propagation to compute the gradient signals. Recently, a class of energy-based models has been studied, which performs equilibrium propagation to compute gradient signals [33, 34, 35, 36, 37]. Focusing on the training dynamics instead of concrete gradient computing, our work is orthogonal to them and is expected to provide insight for algorithm designs of energy-based model training.
2 Analog Training with Generic Response Functions
This section examines the discrete-time dynamics of analog training and introduces the challenges posed by generic response functions. After that, we introduce a family of response functions that reflect crucial physical properties that interest us.
Compact formulations of analog update. We first investigate the dynamics of one element $w$ in $W\in{\mathbb{R}}^{D}$. This paper adopts $w$ to represent the element of the weight $W_{k}$ without specifying its index. The notation makes the formulations more concise and indicates that all elements are updated in parallel. As we discuss in Section˜1, the response granularity $\Delta w_{\min}$ is scaled by the response function $q_{s}(w)$. Naturally, given a desired update $\Delta w$, the increment is approximately scaled by $q_{s}(w)$ as well. Accordingly, we propose that an approximate dynamics of analog update is given by $w^{\prime}\approx U_{q}(w,\Delta w)$, where $U_{q}(w,\Delta w)$ is defined by